Calculus, fall 2013

Study Guide for Chap 02

1. Reading: Sections 2.2-2.6 (up to p104).

2. Table 2.3 in p71 is due to inexact answers of old calculators. It will not happen in modern computing software/hardware. Forget about it.

3. Memorize typical examples of ’limit does not exist’ and discontinuity.

4. Pay attention to variants of lim_{θ→0} sin θ/θ and related examples.

5. Study the precise definitions on page 77, p84, p87, p93 and p104. Find an example for each definition and verify with the ε − δ argument.

6. Study how the ε − δ argument can be used to examine statements like Example 6, 7 in section 2.3 and Theorem 10 in section 2.5.